OpenStudy (anonymous):

find the derivatives of g(x)=x^4(5x-3)^-2

OpenStudy (irishboy123):

\[g(x)={x^4 \over (5x-3)^{2}}\] yes? if so, quitient rule is a very standard approach \[ \left( \dfrac{u}{v}\right)' = \dfrac{u' v - uv'}{v^2}\] \[u = x^4, \qquad v = (5x-3)^2\]

OpenStudy (lfreeman1285):

You could also use the product rule on this one. \[g(x)=x ^{4}(5x-3)^{-2}\] Product rule: \[(uv)\prime=uv \prime+vu\] let's make u=\[x^{4}\] and v=\[(5x+3)^{-2}\] Therefore, you would do \[x ^{4}(-10)(5x+3)+(5x+3)^{-2}4x ^{3}\] After simplifying and combining like terms, and all that stuff, you would have your answer!

OpenStudy (lfreeman1285):

the vu in the product rule is supposed to say \[vu \prime\] sorry if that was confusing